OpenStudy (anonymous):

hi how to appoximate natural logatim without caculator? for example: ln(8/37)?

OpenStudy (gorv):

well \[\ln=\log_{e} \]

OpenStudy (anonymous):

i know that ln in base on e

OpenStudy (gorv):

\[\ln \frac{ 8 }{ 37 }=\log8 +\log37\]

geerky42 (geerky42):

Minus, not plus

OpenStudy (anonymous):

yer minus

OpenStudy (anonymous):

i mean how to approximate answer?

OpenStudy (gorv):

sorry minus

OpenStudy (gorv):

8=2^3

myininaya (myininaya):

2^3=8 2^5=32

myininaya (myininaya):

\[e \approx 2. \text{ blah}\]

OpenStudy (gorv):

ln2=0.69

OpenStudy (anonymous):

ok....in know that: 32<37<64 and ans is between 5*.69 and 6*.69 = 3.45 and 4.14 in mean which number between 3.45 and 4.14 should bee choosen?

OpenStudy (gorv):

no

OpenStudy (gorv):

25<37<49

OpenStudy (gorv):

37 is mean of 25 and 49

OpenStudy (anonymous):

what u say!? if i know 25<37<47 then how i can approximate ln37?

OpenStudy (anonymous):

i only know ln2=.67 and then want to approximate ln 37